OpenStudy (anonymous):

hi can someone help me? ill give medals. A building has x floors. The number of apartments on each floor is 6 less than the number of floors. The number of rooms in each apartment is 1 less than the number of apartments on each floor. The expression below shows the total number of rooms in the building: x(x - 6)(x - 7) ^ that the description i have 3 questions about it.

OpenStudy (anonymous):

What does x(x - 6) represent? thats my first question ^

OpenStudy (anonymous):

x(x-6) represents the no of apartments in building..

OpenStudy (anonymous):

x=no of floors x-6=no of aprtments in each floor.. so x(x-6)=total no of aprtments in x floors.

OpenStudy (anonymous):

What type of polynomial is x(x - 6)(x - 7) second question ^

OpenStudy (anonymous):

total no of rooms in buliding.

OpenStudy (anonymous):

x-6=no of aprtments (x-6)-1=x-7=no of rooms in each aprtment..

OpenStudy (anonymous):

yes but what type of polynomial is it?

OpenStudy (anonymous):

Third degree

OpenStudy (anonymous):

What is the degree of the polynomial x(x - 6)(x - 7) well you just answered my third question ^

OpenStudy (anonymous):

its 3rd degree...

OpenStudy (anonymous):

\[x^3-13 x^2+42 x\]

OpenStudy (anonymous):

max.powers of variable =degree of polynomial

OpenStudy (anonymous):

ok but is it a zero polynomial. monomial Binomial etc

OpenStudy (anonymous):

Trinomal because it has three terms

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

Np